OpenStudy (anonymous):

Write the equation of the line that passes through the points (3,2) and (2,3)

pooja195 (pooja195):

First we need to find the slope

pooja195 (pooja195):

\[\huge~\rm~\frac{ y_2-y_1 }{ x_2-x_1 }=slope\]

OpenStudy (anonymous):

1/-1

pooja195 (pooja195):

Or in other words just -1 what form do they want the equation in?

OpenStudy (anonymous):

y=mx+b

pooja195 (pooja195):

oh ok so now we use the point slope equation \[\huge~\rm~y-y_1=m(x-x_1)\] (x1,y1)=(3,2) m=slope which is -1 just plug in the given information and solve for y

OpenStudy (anonymous):

what would x be?

OpenStudy (texaschic101):

or you can use y = mx + b...either formula will work

pooja195 (pooja195):

x and y are just there they help to figure out the slope and the equation

pooja195 (pooja195):

or you can use @texaschic101 's method

OpenStudy (texaschic101):

and you x and y points.....you can use either set (3,2) or (2,3)

OpenStudy (texaschic101):

which formula do you want to proceed with ?

OpenStudy (anonymous):

y=-3+b... im stuck\

OpenStudy (anonymous):

y=mx+b (3,2)

OpenStudy (texaschic101):

you forgot about your slope(m) of -1

OpenStudy (anonymous):

i did -1(3)

OpenStudy (texaschic101):

y = mx + b slope(m) = -1 (3,2)...x = 3 and y = 2 now we sub 2 = -1(3) + b 2 = -3 + b -- add 3 to both sides 2 + 3 = b 5 = b so your line is : y = -1x + 5 or just y = -x + 5 but if you sub in your other set of points..(2,3) 3 = -1(2) + b 3 = -2 + b 3 + 2 = b 5 = b same thing....y = -x + 5 so either of your points will give you this answer

OpenStudy (texaschic101):

questions ?

OpenStudy (texaschic101):

now if they wanted the answer is point-slope form you would use : y - y1 = m(x - x1). You could also use it for this problem but you would eventually have to change it to y = mx + b form

OpenStudy (anonymous):

no that explains it.

OpenStudy (texaschic101):

glad to help...but please give medal to pooja195....I kinda took over her explaining...should not have done that.